what is the procedure to find the netcapacitance ,charge on each capacitance in complex combination.plz help me sir

Grade:12

1 Answers

Aman Bansal
13 years ago

DEAR GOPALAM VANAPALLI,

IN CASE OF PARALLEL COMBINATION OF CAPACITANCES,

Ceq = e0Aeq/d = e0(A1+A2)/d = e0A1/d + e0A2/d = C1 + C2
Thus, Ceq = C1 + C2.

WHEREAS IN SERIES COMBINATION ,

The quantity that adds is the voltage across each capacitor
DV = DV1 + DV2
The voltage across the capacitors is related to their charge
DV1 = Q/C1   and   DV2 = Q/C2
The definition of the equivalent capacitance is Ceq = Q/DV, or DV = Q/Ceq. Therefore
Q/Ceq = Q/C1 + Q/C2
We can divide out the common factor of Q to arrive at

1/Ceq = 1/C1 + 1/C2
BEST OF LUCK

We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL